OpenStudy (anonymous):

MEDAL TO BEST ANSWER!!!!!!!!! 12x2 - x - 6 Factor this quadratic and show all work

OpenStudy (anonymous):

(3x+2)(4x-3)

OpenStudy (anonymous):

i dont usually show work because i guess and check sorry

OpenStudy (anonymous):

mossyfish try to give the work please

OpenStudy (anonymous):

Formula: \[ax^2+bx_c-a(x-x1)(x-x2)\] So, we have \[12x^2-x-6=12(x-(-2/3))(x-3/4)\] \[12x^2-x-6=12(x+2/3)(x-3/4)\] We can simplify this, let's write 12 as 3*4 Multiply the first bracket by 3 and the second by 4. \[12x^2-x-6=(3x+2)(4x-3)\]

OpenStudy (anonymous):

THANKS!!!! @mossyfish

OpenStudy (anonymous):

You are welcome! :)

OpenStudy (anonymous):

thanx mossyfish i actually didnt know there was a formula. i was taught to basically to guess and check